# Feedback Delay Network Optimization

### v1

@@ Abstract
A common bane of artificial reverberation algorithms is spectral coloration, typically manifesting as metallic ringing, leading to a degradation in the perceived sound quality. This paper presents an optimization framework where a differentiable feedback delay network is used to learn a set of parameters to reduce coloration iteratively. The parameters under optimization include the feedback matrix, as well as the input and output gains. The optimization objective is twofold: to maximize spectral flatness through a spectral loss while maintaining temporal density by penalizing sparseness in the parameter values. A favorable narrower distribution of modal excitation is achieved while maintaining the desired impulse response density. In a subjective assessment, the new method proves effective in reducing perceptual coloration of late reverberation. The proposed method achieves computational savings compared to the baseline while preserving its performance. The effectiveness of this work is demonstrated through two application scenarios where natural-sounding synthetic impulse responses are obtained via the introduction of attenuation filters and an optimizable scattering feedback matrix.
@@ 1 Introduction
Recently, the perceived coloration was attributed to the modal properties of synthesized reverbs, with a wide distribution of modal excitations correlating positively with increased coloration [7]. Therefore, modal decomposition of artificial reverberation architectures, such as FDNs [8], can aid in identifying and mitigating phenomena that are linked to perceptual artifacts. In FDNs, the values of the modal excitations depend on all system parameters [8].
In this paper, the objective is to optimize the gain parameters and feedback matrix of an FDN. The goal is to minimize the perceptual coloration observed in its impulse response (IR). We adopt a stochastic gradient descent approach, focusing specifically on the magnitude of the frequency response. This paper extends our previous work [9] in which part of the loss function was based on the IR. The current work also demonstrates two practical applications of the proposed approach in synthesizing natural-sounding IRs by including attenuation filters in the FDN structure and optimizing the scattering feedback matrix. As we now concentrate solely on the frequency domain, the computational requirements are streamlined. We attain control over temporal density by utilizing its correlation with the density of the feedback matrix entries. Evaluation is conducted by analyzing the modal excitation distribution. A perceptual evaluation against the FDN design before optimization shows that the proposed method successfully decreased perceived coloration.
@@ 2.1 Feedback delay network
Figure 1: FDN with a single input, a single output, and an $N$-by-$N$ feedback matrix A. The thick lines indicate $N$-channel signal paths. The parameters under optimization are highlighted in blue.
An FDN is a recursive system consisting of delay lines, a set of gains, and a scalar feedback matrix through which the delay outputs are coupled to the delay inputs. Fig. 1 presents an example of a single-input, single-output (SISO) FDN architecture. The transfer function of the FDN can be written as
@@ 2.2 Modal decomposition
The IR of the FDN can be expressed as the sum of complex one-pole modes, also known as resonators [8]:
$h(n)=\sum_{i=1}^{\mathcal{{M}}}h_{i}(n)\;.$ (4)
Each mode $h_{i}(n)$ is defined by the pole $\lambda_{i}$ and the residue $\rho_{i}$:
$h_{i}(n)=|\rho_{i}||\lambda_{i}|^{n}e^{\jmath(n\angle\lambda_{i}+\angle\rho_{i})}\,,$ (5)
@@
commonly referred to as the modal decomposition of the FDN [8]. The excitation and initial phase of the $i^{\text{th}}$ mode are determined by the magnitude $|\rho_{i}|$ and phase $\angle\rho_{i}$, respectively, of its corresponding residue. The magnitude and phase of the $i^{\text{th}}$ pole, $|\lambda_{i}|$ and $\angle\lambda_{i}$, respectively, dictate its decay rate and frequency.
A common approach to calculate the modal decomposition (6) from the rational polynomial form of the transfer function (3) involves determining the system poles $\lambda_{i}$. These poles are obtained as the roots of the GCP (2), derived from the eigenvalues of the linearized feedback matrix [8].
@@ 2.3 Homogeneous decay in FDN
Designing an artificial reverberator with FDN often starts by creating a lossless prototype, having an energy-preserving feedback loop [10, 11]. The FDN is said to be lossless if the roots of the GCP (2) have magnitude equal to one, i.e., $|\lambda_{i}|=1$ for all $i$ [12]. The advantage of initially designing a lossless FDN lies in the straightforward implementation of frequency-dependent decay that equally influences all system poles. This is achieved by extending every delay with a frequency-dependent attenuation filter to meet the specified reverberation time [3].
In this study, we mainly focus on the specific case of frequency-independent homogeneous decay. This refers to the case where all modes experience the same decay rate, i.e., $|\lambda_{i}|=\gamma$ for all $i$’s. In Subsection 4.1, we will demonstrate how an FDN optimized at such $\gamma$ can be employed as a lossless prototype for synthesizing reverb with frequency-dependent decay.
Homogeneous decay is achieved with a feedback matrix $\bm{A}$ being the product of a unilossless matrix $\bm{U}$ and a diagonal matrix $\bm{\Gamma}$, whose entries are delay-proportional absorption coefficients, $\bm{\Gamma}=\text{diag}(\gamma^{\bm{m}})$. The feedback matrix can be expressed as
@@
where $f_{\textrm{s}}$ is the sampling rate in Hz and $T_{60}$ is the reverberation time in s defined as the time required for the sound level to decay by $60$ dB from the initial steady-state value.
@@ 2.4 Coloration in FDN
Figure 2: Magnitude response of an FDN with size $N=8$, showcasing the difference between long and short delay lines when the former is 10 times longer than the latter.
Natural late reverberation can be conceptualized as white noise with an exponential decay, providing perceptually ideal smooth diffuse reverberation [14]. In artificial reverberation, achieving a white noise generator involves ensuring the allpass property is maintained. Various perceptual artifacts of recursive reverberators can be attributed, though not exclusively, to deviations from this ideal, which then result in unwanted coloration.
@@
A recent study was conducted to understand further the role of modal excitation in late reverberation coloration [7]. Listening test results suggest that a narrow distribution of the modal excitation values $|\rho_{i}|$ tends to result in a perceptually white spectrum [7]. For large values of $|\rho_{i}|$, coloration starts to become noticeable [7]. We also observed this outcome in our earlier study, where we conducted a listening test focusing on the coloration of FDN-based reverb [9]. In this test, we compared the reverb before and after applying an optimization method to narrow the modal excitation distribution.
The literature also indicates that more than 6000 modes are needed for an IR to be perceived as rather colorless [15]. Additionally, it was shown that for large values of $\mathcal{{M}}$, the modes of the FDNs are uniformly distributed [8], preventing additional coloration that usually results from clusters of modes. Nonetheless, a flat magnitude response and a uniform modal frequency distribution are insufficient to achieve colorlessness.
To comprehend the complexities associated with coloration in FDN, we can examine the comb-filter structure as a special case. Specifically, when the feedback matrix $\bm{A}$ is diagonal, the FDN adopts the configuration of a parallel comb-filter structure. In the case of a homogeneous FDN with a diagonal feedback matrix, the transfer function in (1) is analogous to a superposition of comb filters, with each filter having the transfer function
$H_{i}(z)=\frac{1}{1+\gamma^{m_{i}}z^{-m_{i}}}\,.$ (9)
@@
$\lVert H_{i}(z)\rVert_{2}^{2}$ $=\int_{0}^{2\pi}\lvert H_{i}(e^{\jmath\omega})\rvert^{2}\textrm{d}\omega$ (10)
@@
For any given value of $\gamma$, (11) represents an exponential function. For positive $m_{i}$, the function approaches infinity as $m_{i}$ values decrease and tends to zero as $m_{i}$ values increase. For example, let us consider the case when the gain-per-sample is $\gamma=0.9999$ ($T_{60}=1.44$ s). In this scenario, the energy of the response from the comb filter with $m_{1}=200$ is 8.5 times greater than that with $m_{2}=10m_{1}=2000$. Specifically, $\lVert H_{1}(z)\rVert_{2}^{2}=25.5$, while $\lVert H_{2}(z)\rVert_{2}^{2}=3$. Fundamentally, shorter delays contribute more energy and produce strongly audible metallic-sounding comb peaks. In contrast, longer delays contribute less energy and tend to be masked by the more dominant comb filters [16].
Figure 2 shows a section of the magnitude spectrum of two FDNs with identical parameters but different delay-line lengths. The FDN with shorter delay lengths, approximately one-tenth of the longer delays, exhibits higher magnitude values and more distinct, louder resonances. For long delays, the resonances become less noticeable, and the loudest frequencies exhibit a magnitude approximately 20 dB lower compared to the scenario with shorter delays. This example shows that to achieve a colorless FDN, it is necessary to avoid strongly recirculating short delays and encourage exciting long delays instead.
@@ 2.5 Problem statement
In this paper, our goal is to optimize the feedback delay matrix $\bm{A}$, as well as the input and output gains $\bm{b}$ and $\bm{c}$, to enhance the perceptual colorlessness of the resulting IR. In Fig. 1, the FDN parameters being optimized are indicated in blue. Previous studies indicate that the choice of frequency-dependent attenuation has little impact on coloration [7]. Therefore, we conduct optimization on a long-ringing, frequency-independent prototype FDN. For ideal late reverberator synthesis with this setup, the spectrum’s magnitude must remain constant across frequencies and time. However, the recursive structure of FDN introduces undesired colorations, which constitutes a primary challenge in our work, as emphasized in Subsection 2.4.
@@ 3 FDN optimization
In this section, we present a method to reduce coloration in the IR of an FDN with frequency-independent $T_{60}$. We employ stochastic gradient descent to optimize the parameters of the FDN, which has been made differentiable by utilizing the frequency sampling method.
@@ 3.1 Differentiable FDN
Using the frequency sampling method, the FDN is approximated as an FIR filter. An FIR filter is straightforward to implement, as it does not require a recursive structure. Moreover, an FIR filter can be easily parallelized and optimized for hardware acceleration, such as a graphics processing unit (GPU). To obtain the FIR approximation of $H(z)$, the delay matrix $\bm{D_{m}}(z)$ is evaluated at discrete frequency points
@@
where $M$ indicates the total number of frequency bins equally distributed on the unit circle.
@@
Figure 3: Architecture of the proposed optimization workflow. A thick line indicates a $N$-channel signal path. Backpropagation is highlighted in blue.
@@
Every batch is composed of $\mu<M$ different frequency points that are randomly sampled from $\bm{z}_{M}$. The value of $M$ is such that to ensure oversampling, while we found empirically that $\mu$ can be as small as $0.05f_{\textrm{s}}$. At each epoch, $\bm{H}(z)$ is estimated at all frequency points within the vector $\bm{z}_{M}$. In our prior study [9], the computation of $\bm{H}(z)$ during each training step involved the entire vector $\bm{z}_{M}$ rather than a subset, resulting in a bottleneck due to the computational complexity associated with matrix inversion and multiplication.
@@ 3.2 Orthogonal feedback matrix
To guarantee frequency independence of the energy decay, we adopted the parameterization in (7), where the feedback matrix is computed from an orthogonal matrix $\bm{U}$. The class of orthogonal matrices meets the unitary condition for losslessness and can be obtained through a parameterization of skew-symmetric matrices.
@@ 3.3 Gain-per-sample
Figure 4: Magnitude response of a homogeneous FDN of order $\mathcal{M}=8944$ with random orthogonal feedback matrix and unitary input/output gains at different values of $T_{60}$ in seconds. In the lossless case, $T_{60}=\infty$, the resonances tend to infinity.
@@
We ran optimization at different values of $\gamma$, and we observed good convergence when $T_{60}\leq 10$ s. At inference, $\gamma$ serves as a free parameter, enabling the generation of reverberation with arbitrary $T_{60}$ values.
@@ 3.4 Parameter initialization
In stochastic gradient descent, as in many other optimization algorithms, the choice of initial parameter values can affect the algorithm’s ability to find the global or a good local minimum of the loss function. To start from a point that facilitates an efficient exploration of the loss landscape, we initialize the values of $\bm{b}$ and $\bm{c}$ by drawing them from the normal distribution $\mathcal{N}(0,N^{-1})$, and the entries of $\bm{W}$ from the uniform distribution $\mathcal{U}(-1/\sqrt{N},1/\sqrt{N})$.
The design of the delays is a rather non-trivial task that requires further constraints. To maximize the echo density, the delay lengths should be co-prime [16]. However, the concentration of delays around a certain value may lead to a perceivable strong energy fluctuation over time. Moreover, low-order dependencies, which are integer linear combinations of delays that coincide with other integer linear combinations of delays with small coefficients, can also contribute negatively to the smoothness of the response [6]. To avoid degenerative patterns and ensure a smooth-sounding reverb, we choose delays that are logarithmically distributed co-prime numbers leading to $\mathcal{{M}}\geq 6000$.

### v2

# Efficient Optimization of Feedback Delay Networks for Smooth Reverberation

@@ Abstract
A common bane of artificial reverberation algorithms is spectral coloration, typically manifesting as metallic ringing, leading to a degradation in the perceived sound quality. This paper presents an optimization framework where a differentiable feedback delay network is used to learn a set of parameters to reduce coloration iteratively. The parameters under optimization include the feedback matrix, as well as the input and output gains. The optimization objective is twofold: to maximize spectral flatness through a spectral loss while maintaining temporal density by penalizing sparseness in the parameter values. A favorable narrower distribution of modal excitation is achieved while maintaining the desired impulse response density. In a subjective assessment, the new method proves effective in reducing perceptual coloration of late reverberation. The proposed method achieves computational savings compared to the baseline while preserving its performance. The effectiveness of this work is demonstrated through two application scenarios where smooth-sounding synthetic impulse responses are obtained via the introduction of attenuation filters and an optimizable scattering feedback matrix.
@@ 1 Introduction
Recently, the perceived coloration was attributed to the modal properties of synthesized reverbs, with a wide distribution of modal excitations linked to increased coloration [7]. Therefore, modal decomposition of artificial reverberation architectures, such as FDNs [8], can aid in identifying phenomena that are related to perceptual artifacts. In FDNs, the values of the modal excitations depend on all system parameters [8].
In this paper, the objective is to optimize the gain parameters and feedback matrix of an FDN. Similarly to our previous work [9], the goal is to minimize the perceptual coloration observed in the impulse response (IR) by adjusting the parameters using stochastic gradient descent. In [9], we targeted the flatness of the magnitude response and temporal density by incorporating two loss functions in the frequency and time domains, respectively. For the time domain, the inverse discrete Fourier transform (DFT) operation was needed, which presented a bottleneck that we aim to address in this paper. The current work also demonstrates practical applications of the proposed approach in synthesizing natural-sounding late reverberation by optimizing the scattering feedback matrix and including in the FDN structure attenuation filters that benefit from the optimization. As we now concentrate solely on the frequency domain, the computational requirements are streamlined by analyzing the frequency response sparsely using batch processing over subsets of frequency samples. We attain control over temporal density by utilizing its correlation with the density of the feedback matrix entries instead of computing the system’s IR. Evaluation is conducted by analyzing the modal excitation distribution. A perceptual evaluation against the FDN design before optimization shows that the proposed method successfully decreased perceived coloration.
This work distinguishes itself from a recent and related study presented in [10] by addressing more fundamental issues underlying the limitations in reverb synthesis using FDNs. At the same time, it suggests how this approach can be extended to IR synthesis tasks by introducing a filter design method of our choice. Nonetheless, the main contribution of the presented paper does not depend on the filter used, thereby opening a wider range of applications.
@@ 2.1 Feedback delay network
Figure 1: FDN with a single input, a single output, and an $N$-by-$N$ feedback matrix A. The thick lines indicate $N$-channel signal paths. The circled multiplication symbols indicate a vector multiplication. The parameters under optimization are highlighted in blue.
A FDN is a recursive system consisting of delay lines, a set of gains, and a scalar feedback matrix through which the delay outputs are coupled to the delay inputs. Fig. 1 presents an example of a single-input, single-output (SISO) FDN architecture. The transfer function of the FDN can be written as
@@ 2.2 Modal decomposition
The IR of the FDN can be expressed as the sum of complex one-pole modes, also known as resonators, each defined by its pole $\lambda_{i}$ and its residue $\rho_{i}$ [8]:
$h(n)=\sum_{i=1}^{\mathcal{{M}}}|\rho_{i}||\lambda_{i}|^{n}e^{\jmath(n\angle\lambda_{i}+\angle\rho_{i})}\;.$
@@
commonly referred to as the modal decomposition of the FDN [8]. The excitation and initial phase of the $i^{\text{th}}$ mode are determined by the magnitude $|\rho_{i}|$ and phase $\angle\rho_{i}$, respectively, of its corresponding residue. The magnitude and phase of the $i^{\text{th}}$ pole, $|\lambda_{i}|$ and $\angle\lambda_{i}$, respectively, dictate its decay rate and frequency. A common approach to calculate the modal decomposition (4) from the rational polynomial form of the transfer function (3) involves determining the system poles $\lambda_{i}$ from the eigenvalues of the linearized feedback matrix [8].
An important aspect for the remainder of this paper is the distribution of the modal excitation values $|\rho_{i}|$. While the decay of the modes $|\lambda_{i}|$ is typically governed entirely by the target reverberation time $T_{60}$, the modal excitation remains largely unconstrained by design. The connection between the distribution of modal excitation with perceived coloration [7] is a key finding underlying the proposed optimization method
@@ 2.3 Homogeneous decay in FDN
Designing an artificial reverberator with FDN often starts by creating a lossless prototype, having an energy-preserving feedback loop [11, 12]. The FDN is said to be lossless if the roots of the GCP (2) have magnitude equal to one, i.e., $|\lambda_{i}|=1$ for all $i\,$s [13]. The advantage of initially designing a lossless FDN lies in the straightforward implementation of frequency-dependent decay that equally influences all system poles. This is achieved by extending every delay with a frequency-dependent attenuation filter to meet the specified reverberation time [3].
In this study, we mainly focus on the specific case of frequency-independent homogeneous decay. This refers to the case where all modes experience the same decay rate, i.e., $|\lambda_{i}|=\gamma$ for all $i$ s. In Subsection 4.1, we will demonstrate how an FDN optimized as such can be employed as a lossless prototype for synthesizing reverb with frequency-dependent decay.
Homogeneous decay is achieved with a feedback matrix $\bm{A}$ being the product of a unilossless matrix $\bm{U}$ and a diagonal matrix $\bm{\Gamma}$, whose entries are delay-proportional absorption coefficients, $\bm{\Gamma}=\text{diag}(\gamma^{\bm{m}})$, $\gamma^{\bm{m}}$ represents the $\gamma$ value raised to the power of each corresponding delay line in $\bm{m}$. The feedback matrix can be expressed as
@@
where $f_{\textrm{s}}$ is the sampling rate in Hz.
@@ 2.4 Coloration in FDN
Figure 2: Magnitude response of a homogeneous FDN with size $N=8$, diagonal matrix, and unit input and output gains, showcasing the difference between long and short delay lines when the former is 10 times longer than the latter.
Natural late reverberation can be modeled as white noise with an exponential decay, providing perceptually ideal smooth diffuse reverberation [15]. In artificial reverberation, achieving a white noise generator involves ensuring the allpass property is maintained. Various perceptual artifacts of recursive reverberators can be attributed, though not exclusively, to deviations from this ideal, which then result in unwanted coloration.
@@
A recent study was conducted to understand further the role of modal excitation in late reverberation coloration [7]. Listening test results suggest that a narrow distribution of the modal excitation values $|\rho_{i}|$ tends to result in a perceptually white spectrum [7]. However, when a subset of modes exhibit high $|\rho_{i}|$ values relative to the mean of the system’s modal excitation distribution, coloration starts to become noticeable [7]. We also observed this outcome in our earlier study, where we conducted a listening test focusing on the coloration of a FDN-based reverb [9]. In this test, we compared the reverb before and after applying an optimization method, which resulted in a narrower modal excitation distribution.
The literature indicates that more than 6000 modes are needed for an IR to be perceived as rather colorless [16]. Additionally, it was shown that for large values of $\mathcal{{M}}$, the modes of the FDNs are uniformly distributed [8], preventing additional coloration that usually results from clusters of modes. Nonetheless, a flat magnitude response and a uniform modal frequency distribution are insufficient to achieve colorlessness.
To comprehend the complexities associated with coloration in FDN, we can examine the comb-filter structure as a special case. Specifically, when the feedback matrix $\bm{A}$ is diagonal, the FDN adopts the configuration of a parallel comb-filter structure. In the case of a homogeneous FDN with a diagonal feedback matrix, the transfer function in (1) is analogous to a superposition of comb filters with the delay line in the feedforward path and the gain $\gamma^{m}_{i}$ in the feedback. Each filter has a transfer function
$H_{\textrm{comb}_{i}}(z)=\frac{z^{-m_{i}}}{1-\gamma^{m_{i}}z^{-m_{i}}}\,.$ (7)
@@
$\lVert H_{\textrm{comb}_{i}}(z)\rVert_{2}^{2}$ $=\int_{0}^{2\pi}\lvert H_{\textrm{comb}_{i}}(e^{\jmath\omega})\rvert^{2}\textrm{d}\omega$ (8)
@@
For any given value of $\gamma$, (9) represents an exponential function. For positive $m_{i}$, the function approaches infinity as $m_{i}$ values decrease and tends to zero as $m_{i}$ values increase. For example, let us consider the case when the gain-per-sample is $\gamma=0.9999$ ($T_{60}=1.44$ s). In this scenario, the energy of the response from the comb filter with $m_{1}=200$ is 8.5 times greater than that with $m_{2}=10m_{1}=2000$. Specifically, $\lVert H_{1}(z)\rVert_{2}^{2}=25.5$, while $\lVert H_{2}(z)\rVert_{2}^{2}=3$. Fundamentally, shorter delays contribute more energy and produce strongly audible metallic-sounding comb peaks. In contrast, longer delays contribute less energy and tend to be masked by the more dominant comb filters [17]. The same conclusion can be derived through pole-zero analysis of (7). By definition, FDNs are networks of comb filters whose feedback paths are interconnected through a feedback matrix. Therefore, the reasoning above applies to FDNs with dense feedback matrices without loss of generality.
Figure 2 shows a section of the magnitude spectrum of two FDNs with identical parameters but different delay-line lengths. Since the resonant frequencies are uniformly distributed across the spectrum, [8] the choice of the observed frequency range is arbitrary and serves only for visualization purposes. The feedback matrix is diagonal with values alternating between +1 and -1. The input and output gains are unit vectors. The FDN with shorter delay lengths, approximately one-tenth of the longer delays, exhibits higher magnitude values and more distinct, louder resonances. For long delays, the resonances become less noticeable, and the loudest frequencies exhibit a magnitude approximately 20 dB lower compared to the scenario with shorter delays. This example suggests that to achieve a colorless FDN, it is necessary to avoid strongly recirculating short delays and instead focus on exciting long delays.
@@ 2.5 Problem statement
In this paper, our goal is to optimize the feedback delay matrix $\bm{A}$, as well as the input and output gains $\bm{b}$ and $\bm{c}$, to enhance the perceptual colorlessness of the resulting IR. In Fig. 1, the FDN parameters being optimized are indicated in blue. Previous studies indicate that the choice of frequency-dependent attenuation has little impact on coloration [7] Therefore, we conduct optimization on a long-ringing, frequency-independent prototype FDN. For ideal late reverberator synthesis with this setup, the spectrum’s magnitude must remain constant across frequency and time. However, the recursive structure of FDN introduces undesired colorations, which constitutes a primary challenge in our work, as emphasized in Subsection 2.4.
@@ 3 FDN optimization
In this section, we present a method to reduce coloration in the IR of an FDN with frequency-independent $T_{60}$. We employ stochastic gradient descent to optimize the gain parameters of the FDN, which has been made differentiable by utilizing the frequency sampling method.
@@ 3.1 Differentiable FDN
Using the frequency sampling method, the FDN is approximated as an FIR filter. An FIR filter is straightforward to implement, as it does not require a recursive structure. Moreover, an FIR filter can be easily parallelized and optimized for hardware acceleration, such as a graphics processing unit. To obtain the FIR approximation of $H(z)$, the delay matrix $\bm{D_{m}}(z)$ is evaluated at discrete frequency points
@@
where $M$ indicates the total number of frequency bins evenly distributed on the unit circle.
@@
Figure 3: Architecture of the proposed optimization workflow. A thick line indicates an $N$-channel signal path. Backpropagation is highlighted with a dashed blue line.
@@
Every batch is composed of $\mu<M$ different frequency points that are randomly sampled from $\bm{z}_{M}$. The value of $M$ is such that to ensure oversampling, we found empirically that $\mu$ can be as small as $0.05f_{\textrm{s}}$. At each epoch, $\bm{H}(z)$ is estimated at all frequency points within the vector $\bm{z}_{M}$. In our prior study [9], the computation of $\bm{H}(z)$ during each training step involved the entire vector $\bm{z}_{M}$ rather than a subset, resulting in a bottleneck due to the computational complexity associated with matrix inversion and multiplication.
@@ 3.2 Orthogonal feedback matrix
To guarantee frequency independence of the energy decay, we adopted the parameterization in (5), where the feedback matrix is computed from an orthogonal matrix $\bm{U}$. The class of orthogonal matrices meets the unitary condition for losslessness and can be obtained through the parameterization of skew-symmetric matrices described in [18].
@@ 3.3 Gain-per-sample
Figure 4: Magnitude response of a homogeneous FDN of size $N=8$ and order $\mathcal{M}=8944$ with random orthogonal feedback matrix and unitary input/output gains at different values of $T_{60}$ in seconds. In the lossless case, $T_{60}=\infty$, the resonances tend to infinity.
@@
While (6) provides a lower bound for $\gamma$, defining an upper bound is more complex. Therefore, we recommend treating it as a hyperparameter to be tuned for optimal performance. We ran optimization at different values of $\gamma$, and we observed good convergence when $T_{60}\leq 10$ s. At inference, $\gamma$ serves as a free parameter, enabling the generation of reverberation with arbitrary $T_{60}$ values.
@@ 3.4 Parameter initialization
In stochastic gradient descent, as in many other optimization algorithms, the choice of initial parameter values can affect the algorithm’s ability to find the global or a good local minimum of the loss function. We initialize the values of $\bm{b}$ and $\bm{c}$ by drawing them from the normal distribution $\mathcal{N}(0,N^{-1})$, and the entries of $\bm{W}$ from the uniform distribution $\mathcal{U}(-1/\sqrt{N},1/\sqrt{N})$. This initialization strategy has empirically yielded good results in our experiments.
The design of the delays is a rather non-trivial task that requires further constraints. To maximize the echo density, the delay lengths should be co-prime [17]. However, the concentration of delays around a certain value may lead to a perceivable strong energy fluctuation over time. Moreover, low-order dependencies, which are integer linear combinations of delays that coincide with other integer linear combinations of delays with small coefficients, can also contribute negatively to the smoothness of the response [6]. To avoid degenerative patterns and ensure a smooth-sounding reverb, we choose delays that are logarithmically distributed co-prime numbers leading to $\mathcal{{M}}\geq 6000$. The design of delays depends on the specific application, and the gain parameters must be optimized whenever a new set of delay lines is used since they contribute to the system’s modal excitation distribution.